OpenStudy (anonymous):

differentiate f(x) = e^sqrt(3x)

OpenStudy (freckles):

\[f'(x)=(\sqrt{3x})'e^{\sqrt{3x}}\]

OpenStudy (anonymous):

Is there any way you can help me out step by step?

OpenStudy (freckles):

So you do not know how to differentiate sqrt(3x)?

OpenStudy (freckles):

Do you know how to differentiate sqrt(x)?

OpenStudy (anonymous):

What I have so far is f '(x) = (e^(3x))(sqrt(3x))'

OpenStudy (anonymous):

differentiating sqrt(3x) is where I am stuck

OpenStudy (freckles):

Well do you know to differentiate sqrt(x)?

OpenStudy (freckles):

\[\sqrt{3x}=\sqrt{3} \sqrt{x}\]

OpenStudy (freckles):

\[\frac{d}{dx} \sqrt{3x}=\sqrt{3} \frac{d}{dx} \sqrt{x}\]

OpenStudy (freckles):

recall sqrt(x) is x^(1/2)